Mard Manufacturing has 4,144,282 shares of common stock outstanding. The current share price is $50.23, and the book value per share is $7.6. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $42,105,520, has a 0.08 coupon, matures in 28 years and sells for 96 percent of par. The second issue has a face value of $75,557,201, has a 0.05 coupon, matures in 11 years, and sells for 88 percent of par. What is Mard weight of debt on a market value basis? (4 decimals )
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=50.23*4144282
=208167284.86
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=42105520*1*0.96
=40421299.2
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=75557201*1*0.88
=66490336.88
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=208167284.86+40421299.2+66490336.88
=315078920.94
Weight of debt = MV of Bond/MV of firm
Weight of debt = 106911636.08/315078920.94
W(D)=0.3393